OpenStudy (anonymous):

Use linear Approximation to approximate e^.01

OpenStudy (anonymous):

ah a relic from the days when you could just whip out a $8 calculator and plug it in

OpenStudy (anonymous):

which, btw, is exactly what i would do!

myininaya (myininaya):

f(x)=e^x we need to find a tangent line close to x=0.01 x=0 is pretty close f(0)=e^0=1 <-(0,1) oour point of tangency f'(x)=e^x f'(0)=e^0=1 <- this is slope=m y=mx+b y=1x+b we know a point on the line (0,1) 1=1*0+b 1=b y=1x+1 or y=x+1 so we can estimate e^.01 by pluging 0.01 into our linear approximation f(x) apprx=0.01+1=1.01